# A Survey of World Models for Autonomous Driving

### v3

@@
Abstract
Recent breakthroughs in autonomous driving have been propelled by advances in robust world modeling, fundamentally transforming how vehicles interpret dynamic scenes and execute safe decision-making. World models have emerged as a linchpin technology, offering high-fidelity representations of the driving environment that integrate multi-sensor data, semantic cues, and temporal dynamics. This paper systematically reviews recent advances in world models for autonomous driving, proposing a three-tiered taxonomy: (i) Generation of Future Physical World, covering Image-, BEV-, OG-, and PC-based generation methods that enhance scene evolution modeling through diffusion models and 4D occupancy forecasting; (ii) Behavior Planning for Intelligent Agents, combining rule-driven and learning-based paradigms with cost map optimization and reinforcement learning for trajectory generation in complex traffic conditions; (ii) Interaction between Prediction and Planning, achieving multi-agent collaborative decision-making through latent space diffusion and memory-augmented architectures. The study further analyzes training paradigms, including self-supervised learning, multimodal pretraining, and generative data augmentation, while evaluating world models’ performance in scene understanding and motion prediction tasks. Future research must address key challenges in self-supervised representation learning, long-tail scenario generation, and multimodal fusion to advance the practical deployment of world models in complex urban environments. Overall, the comprehensive analysis provides a technical roadmap for harnessing the transformative potential of world models in advancing safe and reliable autonomous driving solutions.
Index Terms:
@@ Abstract
1 Introduction
The quest for autonomous driving is a focal point in both scientific research and industry endeavors. It aims to reduce traffic accidents, alleviate congestion, and enhance mobility for societal groups [1]. Current statistics underscore that human error remains the principal cause of accidents [2], indicating that reducing reliance on direct human control could lower the incidence of traffic-related fatalities and injuries. Beyond safety, economic factors also propel the development of autonomous driving technologies [3].
Despite these incentives, achieving high-level autonomy faces major hurdles. Foremost among these is perceiving and understanding dynamic traffic scenarios, which requires fusing heterogeneous sensor streams into an environmental representation [4, 5]. Autonomous vehicles must rapidly assimilate multi-modal data, detect salient objects, and anticipate their motion under adversarial conditions – such as sensor degradation in heavy rain, unpaved roads with poor markings, or aggressive traffic maneuvers [6, 7]. Moreover, real-time decision-making introduces computational constraints, imposing millisecond-level responsiveness to address unexpected obstacles or anomalous behaviors [8]. Equally pivotal is the system’s resilience in extreme or long-tail scenarios (e.g., severe weather, construction zones, or erratic driving behaviors), where performance shortfalls can compromise overall safety [9, 10].
Within this context, constructing robust world models has emerged as a cornerstone. In essence, a world model is a generative spatio-temporal neural system that compresses multi-sensor physical observations into a compact latent state and rolls it forward under hypothetical actions, letting the vehicle rehearse futures before they occur [11, 12]. This high-fidelity internal map [3] updates online, and feeds downstream tasks such as physical-world prediction [13]. Recent work refines such models with generative sensor fusion [14, 15, 16, 17, 18], unifying heterogeneous inputs into consistent top-down views [19, 20].
These robust world models leverage environmental representations to optimize behavior planning, serving as a keystone for safer, efficient autonomous driving. By enabling trajectory optimization, real-time hazard detection, and adaptive route planning, they reduce predictable risks [5] and are compatible with vehicle-to-everything systems [8]. World models facilitate cohesive integration between perception and control subsystems, streamlining the closed-loop autonomy pipeline [21, 22].
@@ 1 Introduction
Figure 1: Structure of the overall review1). The top row outlines the organization. The second and third rows illustrate the background and key components. The fourth row highlights various methodologies for training models in autonomous driving. The bottom row showcases four application areas for world models in autonomous driving.
Differences between Our Survey and Others. Existing surveys on world models for autonomous driving can be classified into two categories. The mainstream category covers general world models used across many fields [23, 11], with autonomous driving only one area. The second category [24, 25], though relatively scarce, focuses on the application of world models within autonomous driving. They categorize studies coarsely and often focus solely on world simulation or lack discussions on the interaction between planning and prediction, resulting in a lack of a clear taxonomy. ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ sector.
@@
Current research falls into three key areas: (i) Generation of Future Physical World: Focusing on the physical world evolution of both dynamic objects and static entities [10, 26, 18]; (ii) Behavior Planning for Intelligent Agents: Examining generative and rule-based planning methods that produce safe, efficient paths under uncertain driving conditions [13, 27, 28]; (iii) Interaction between Behavior Planning and Future Prediction: Highlighting how unified frameworks can capture agent interactions and leverage predictive insights for collaborative optimization [29, 29]. Specifically, the following contributions are presented:
[…]
    ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ ${}_{\!}$ generation.
[…]
@@
[…]
    Proposition of Interactive Model Research: Interactive models that jointly address future prediction and agent behavior are reviewed, indicating how this synergy can enhance real-world adaptability and operational safety.
@@
Figure 2: The paradigms of generation methods for future physical world3.1). (a) Image-based generation synthesizes high-fidelity 2D images. (b) BEV-based generation forecasts BEV maps using paired image and BEV cues. (c) OG-based generation predicts 4D occupancy grids. (d) PC-based generation outputs future LiDAR sweeps. Boxes indicate processing modules; solid arrows denote data flow.
2 Background
@@
2.1 Definition of World Models
A world model is a generative spatio-temporal neural system that encodes the external physical environment into a compact latent state, jointly capturing geometry, semantics and causal context [11]. This internal state is learned without labels: the system first employs a self-learning compressor to squeeze raw sensor frames into a handful of key numbers; next, a time-aware prediction module uses the hidden state and the agent’s action to infer what the next hidden state will be, allowing the agent to rehearse an entire trajectory in its “mind” before acting in the real world [30, 12]. Because every part of the pipeline is differentiable [12], the whole model functions as an interpretable virtual sandbox that supplies gradients, data samples, and “what-if” roll-outs, enabling the agent to make decisions more efficiently [11].
World Models for Autonomous Driving apply the same idea to road domains by mapping synchronized camera images, LiDAR sweeps, radar echoes and HD-maps into a single latent scene graph, thereby unifying perception and prediction within one representation [24]. Inside this latent space the model rolls forward the joint dynamics of surrounding traffic and the ego vehicle over long horizons [11], enabling realistic scene generation, multi-agent behaviour forecasting, trajectory evaluation and low-latency control – all while respecting the physical constraints and multimodal uncertainties endemic to real-world driving.
2.2 Core Tasks in World Models for Autonomous Driving
@@
$\!\!\!\!\bm{z}^{T+1},\!\tau^{T+1}\!=\!\bm{w}((\bm{I}^{T},\!\cdots,\!\bm{I}^{T-t}),\!(\bm{P}^{T},\!\cdots,\!\bm{P}^{T-t})).$ (1)
The first core task is generation of future physical world [31, 32, 20, 33], which involves forecasting the future states of dynamic entities. This task emphasizes capturing potential interactions, stochastic behaviors, and uncertainties within rapidly changing and complex scenes. The second core task is behavior planning for intelligent agents [27, 34, 22, 35], focusing on generating optimal and feasible trajectories for the ego vehicle. It requires accounting for safety constraints, dynamic obstacles, traffic regulations, and real-time adaptability.
3 Taxonomy
3.1 Generation of Future Physical World
@@
TABLE I: Overview of summary, methods and core architectures for Image-/BEV-based generation (§3.1.1 and §3.1.2).
@@
Image-based Generation Dreamer Series DriveDreamer[10] ECCV’24 Diffusion Model Text + Image + HDMap + Box + Actions 2D Image + Actions nuScenes[36]
@@
DriveDreamer-2[37] AAAI’25 LLM + Diffusion Model Text + HDMap + Box 2D Image nuScenes[36]
@@
DriveDreamer4D[26] CVPR’24 Diffusion Model + 4DGS Text + Image + Trajectory 2D Image Waymo[38]
ReconDreamer[39] CVPR’24 Diffusion Model Image + HDMap + Box 2D Image Waymo[38]
@@
WorldDreamer[40] arXiv’24 LLM Image + Video + Text + Action 2D Image nuScenes[36]
@@
CarDreamer[41] IOTJ’24 World Model Backbone Image + BEV + LiDAR + Radar 2D Image + Actions CARLA[42]
[…]
DrivingDiffusion[14] ECCV’24 Diffusion Model Image + Layouts + Text + Optical Flow 2D Image CARLA[42]
Drive-WM[44] CVPR’24 Diffusion Model Image + Action + Box + Map + Text 2D Image nuScenes[36] + Waymo[38]
DriVerse [45] CVPR’25 Diffusion Model Image + Trajectory 2D Image nuScenes[36] + Waymo[38]
GEM[16] CVPR’24 Diffusion Model Image + Trajectory + Human pose + DINO features 2D Image + Depth Self-created dataset[16]
Vista[46] NeurIPS’24 Diffusion Model Image + Action 2D Image Self-created dataset[46]
SimGen [47] NeurIPS’24 Diffusion Model Image + Depth + Segmentation + Text 2D Image DIVA [47] + nuScenes[36]
Delphi[48] arXiv’24 Diffusion Model Image + BEV layout + Failure Case 2D Image nuScenes[36]
BevWorld[32] arXiv’24 Diffusion Model Image + Point cloud + Action 2D Image + 3D LiDAR nuScenes[36] + CARLA[49]
DrivePhysica [50] arXiv’24 Diffusion Model Image + Poses + Box + Lane + Text 2D Image nuScenes[36]
Imagine-2-Drive [51] arXiv’24 Diffusion Model Image 2D Image CARLA[49]
MaskGWM[52] arXiv’25 Diffusion Model Image + Text + Action 2D Image OpenDV-2K[53] + nuScenes[36]
@@
EOTWM [54] arXiv’25 Diffusion Model Image + Text + Trajectory 2D Image nuScenes[36]
UniFuture [55] arXiv’25 Diffusion Model Image + Depth 2D Image + Depth nuScenes[36]
DiVE [56] arXiv’25 Diffusion Model Image + Text + Box + Road condition 2D Image nuScenes[36]
CoGen [57] arXiv’25 Diffusion Model Image + Occupancy 2D Image nuScenes[36]
GAIA-2 [58] arXiv’25 Diffusion Model Image + Action + Text 2D Image Internal dataset[58]
MiLA [59] arXiv’25 Diffusion Model Image + Waypoints + Text + Camera parameters 2D Image nuScenes[36]
@@
DiST-4D [60] arXiv’25 Diffusion Model Image + Trajectory + BEV + Pose + Box RGBD image nuScenes[36] + Waymo[38]
[…]
HoloDrive[61] arXiv’24 Transformer Image + LiDAR + Layouts + Text 2D Image + 3D LiDAR nuScenes[36]
BEVGen[62] RA-L’24 Transformer Image + BEV layout 2D Image nuScenes[36] + Argoverse 2[63]
DrivingWorld[64] arXiv’24 Video GPT Image + Orientation + Location 2D Image NuPlan[65] Internal data[64]
@@
DriveSim [66] arXiv’24 MLLMs Image + Text 2D Image Internal dataset[66]
@@
TMPE [35] arXiv’24 Tansformer Image + Path + Speed 2D Image + BEV CARLA [42]
@@
NeMo[67] ECCV’24 Transformer Image 2D Image + LiDAR + Occupancy nuScenes[36]
@@
FUTURIS[68] arXiv’25 Transformer Segmentation + Depth Segmentation + Depth Cityscapes[69]
@@
BEV-based Generation FIERY [33] ICCV’21 Transformer Image + Ego motion BEV nuScenes[36]
StretchBEV[70] ECCV’22 LSTM Image BEV nuScenes[36]
MILE[34] NeurIPS’22 RNN Image + BEV segmentation + Action Image + BEV + Action CARLA[49]
UNO[31] CVPR’24 CNN LiDAR Occupancy + LiDAR + BEV nuScenes[36]
PowerBEV [71] IJCAI’23 CNN Image BEV Instance nuScenes[36]

### v4

@@
Abstract.
Recent breakthroughs in autonomous driving have been propelled by advances in robust world modeling, fundamentally transforming how vehicles interpret dynamic scenes and execute safe decision-making. World models have emerged as a linchpin technology, offering high-fidelity representations of the driving environment that integrate multi-sensor data, semantic cues, and temporal dynamics. This paper systematically reviews recent advances in world models for autonomous driving, proposing a three-tiered taxonomy: (i) Generation of Future Physical World, covering Image-, BEV-, OG-, and PC-based generation methods that enhance scene evolution modeling through diffusion models and 4D occupancy forecasting; (ii) Behavior Planning for Intelligent Agents, combining rule-driven and learning-based paradigms with cost map optimization and reinforcement learning for trajectory generation in complex traffic conditions; (ii) Interaction between Prediction and Planning, achieving multi-agent collaborative decision-making through latent space diffusion and memory-augmented architectures. The study further analyzes training paradigms, including self-supervised learning, multimodal pretraining, and generative data augmentation, while evaluating world models’ performance in scene understanding and motion prediction tasks. Future research must address key challenges in self-supervised representation learning, multimodal fusion, and advanced simulation to advance the practical deployment of world models in complex urban environments. Overall, the comprehensive analysis provides a technical roadmap for harnessing the transformative potential of world models in advancing safe and reliable autonomous driving solutions.
@@ Abstract.
$c$s: General and reference Surveys and overviewsccs: Computing methodologies Computer visionccs: Networks Network design and planning algorithms
1. Introduction
The quest for autonomous driving has become a focal point in both scientific research and industry endeavors. At its core lies the ambition to reduce traffic accidents, alleviate congestion, and enhance mobility for societal groups (Fagnant and Kockelman, 2015). Current statistics underscore that human error remains the principal cause of accidents (Organization, 2019), indicating that reducing reliance on direct human control could lower the incidence of traffic-related fatalities and injuries. Beyond safety, economic factors also propel the development of autonomous driving technologies (Liu et al., 2021).
Despite these incentives, achieving high-level autonomy faces major hurdles. Foremost among these is perceiving and understanding dynamic traffic scenarios, which requires fusing heterogeneous sensor streams into an environmental representation (Grigorescu et al., 2020; Furda and Vlacic, 2011). Autonomous vehicles must rapidly assimilate multi-modal data, detect salient objects, and anticipate their motion under adversarial conditions – such as sensor degradation in heavy rain, unpaved roads with poor markings, or aggressive traffic maneuvers (Schwarting et al., 2018; Manivasagam et al., 2020). Moreover, real-time decision-making introduces computational constraints, imposing millisecond-level responsiveness to address unexpected obstacles or anomalous behaviors (Li, 2023). Equally pivotal is the system’s resilience in extreme or long-tail scenarios (e.g., severe weather, construction zones, or erratic driving behaviors), where performance shortfalls can compromise overall safety (Bogdoll et al., 2024; Wang et al., 2024c).
Within this context, constructing robust world models has emerged as a cornerstone. In essence, a world model is a generative spatio-temporal neural system that compresses multi-sensor physical observations into a compact latent state and rolls it forward under hypothetical actions, letting the vehicle rehearse futures before they occur (Ding et al., 2024b; Ha and Schmidhuber, 2018b). This high-fidelity internal map (Liu et al., 2021) updates online, and feeds downstream tasks such as physical-world prediction (Qian et al., 2019). Recent work refines such models with generative sensor fusion (Li et al., 2024d; Hu et al., 2023a; Hassan et al., 2024; Bian et al., 2025; Wang et al., 2024b), unifying heterogeneous inputs into consistent top-down views (Wang et al., 2020; Yang et al., 2024a).
These robust world models leverage environmental representations to optimize behavior planning, serving as a keystone for safer, efficient autonomous driving. By enabling trajectory optimization, real-time hazard detection, and adaptive route planning, they reduce predictable risks (Furda and Vlacic, 2011) and are compatible with vehicle-to-everything systems (Li, 2023). World models facilitate cohesive integration between perception and control subsystems, streamlining the closed-loop autonomy pipeline (Min et al., 2024b; Zheng et al., 2025b).
@@ 1. Introduction
Figure 1. Structure of the overall review1). The top row outlines the organization. The second and third rows illustrate the background and key components. The fourth row highlights various methodologies for training models in autonomous driving. The bottom row showcases four application areas for world models in autonomous driving.
Differences between Our Survey and Others. Existing surveys on world models for autonomous driving can be classified into two categories. The mainstream category covers general world models used across many fields (Zhu et al., 2024; Ding et al., 2024b), with autonomous driving only one area. The second category (Guan et al., 2024; Fu et al., 2024b), though relatively scarce, focuses on the application of world models within autonomous driving. They categorize studies coarsely and often focus solely on world simulation or lack discussions on the interaction between planning and prediction, resulting in a lack of a clear taxonomy. In this paper, a comprehensive review of world models for autonomous driving is presented, and their applications are explored, thereby highlighting how these models are shaped by and adapt to the automotive sector.
@@
Current research falls into three key areas: (i) Generation of Future Physical World: Focusing on the physical world evolution of both dynamic objects and static entities (Wang et al., 2024c; Zhao et al., 2025a; Wang et al., 2024b); (ii) Behavior Planning for Intelligent Agents: Examining generative and rule-based planning methods that produce safe, efficient paths under uncertain driving conditions (Qian et al., 2019; Hu et al., 2022a; Casas et al., 2021); (iii) Interaction between Behavior Planning and Future Prediction: Highlighting how unified frameworks can capture agent interactions and leverage predictive insights for collaborative optimization (Yang et al., 2024d, d). Specifically, the following contributions are presented:
  • Analysis of Future Prediction Models: Image-/BEV-/OG-/PC-based generation methods are examined to show how geometric and semantic fidelity is achieved, including 4D occupancy forecasting and diffusion-based generation.
@@
  • Proposition of Interactive Model Research: Interactive models that jointly address future prediction and agent behavior are systematically reviewed, indicating how this synergy can vastly enhance real-world adaptability and operational safety.
@@
2. Background
@@
2.1. Definition of World Models
A world model is a generative spatio-temporal neural system that encodes the external physical environment into a compact latent state, jointly capturing geometry, semantics and causal context (Ding et al., 2024b). This internal state is learned without labels: the system first employs a self-learning compressor to squeeze raw sensor frames into a handful of key numbers; next, a time-aware prediction module uses the hidden state and the agent’s action to infer what the next hidden state will be, allowing the agent to rehearse an entire trajectory in its “mind” before acting in the real world (Ha and Schmidhuber, 2018a, b). Because every part of the pipeline is differentiable (Ha and Schmidhuber, 2018b), the whole model functions as an interpretable virtual sandbox that supplies gradients, data samples, and “what-if” roll-outs, enabling the agent to make decisions more efficiently (Ding et al., 2024b).
World Models for Autonomous Driving apply the same idea to road domains by mapping synchronized camera images, LiDAR sweeps, radar echoes and HD-maps into a single latent scene graph, thereby unifying perception and prediction within one representation (Guan et al., 2024). Inside this latent space the model rolls forward the joint dynamics of surrounding traffic and the ego vehicle over long horizons (Ding et al., 2024b), enabling realistic scene generation, multi-agent behaviour forecasting, trajectory evaluation and low-latency control – all while respecting the physical constraints and multimodal uncertainties endemic to real-world driving.
2.2. Core Tasks in World Models for Autonomous Driving
@@
(1) $\!\!\!\!\bm{z}^{T+1},\!\tau^{T+1}\!=\!\bm{w}((\bm{I}^{T},\!\cdots,\!\bm{I}^{T-t}),\!(\bm{P}^{T},\!\cdots,\!\bm{P}^{T-t})).$
The first core task is generation of future physical world (Agro et al., 2024; Zhang et al., 2024a; Yang et al., 2024a; Hu et al., 2021b), which involves forecasting the future states of dynamic entities. This task emphasizes capturing potential interactions, stochastic behaviors, and uncertainties within rapidly changing and complex scenes. The second core task is behavior planning for intelligent agents (Hu et al., 2022a, b; Zheng et al., 2025b; Popov et al., 2024), focusing on generating optimal and feasible trajectories for the ego vehicle. It requires accounting for safety constraints, dynamic obstacles, traffic regulations, and real-time adaptability.
3. Taxonomy
3.1. Generation of Future Physical World
@@
Figure 2. The paradigms of generation methods for future physical world3.1). (a) Image-based generation synthesizes high-fidelity 2D images. (b) BEV-based generation forecasts BEV maps using paired image and BEV cues. (c) OG-based generation predicts 4D occupancy grids. (d) PC-based generation outputs future LiDAR sweeps. Boxes indicate processing modules; solid arrows denote data flow.
3.1.1. Image-based Generation
Dreamer Series. Dreamer series was first proposed for learning world models to address dynamic scene understanding in gaming and robotics (Hafner et al., 2020, 2021, 2023), later expanded to autonomous driving with DriveDreamer (Wang et al., 2024c). It enables controllable video generation in high-density urban traffic using real-world driving data, though they remain limited to 2D outputs lacking spatial-temporal coherence. Subsequent research introduces DriveDreamer-2 (Zhao et al., 2025b), incorporating LLM-based prompting mechanisms to enhance interactivity and diversity. With further advancements, DriveDreamer4D (Zhao et al., 2025a) significantly extends model capabilities by utilizing world model priors to generate spatial-temporally coherent 4D driving videos, marking a transition from traditional 2D to advanced 4D video generation. ReconDreamer (Ni et al., 2024) explores online restoration techniques to achieve precise video reconstruction of dynamic scenes. WorldDreamer (Wang et al., 2024d) expands the applicability of world models by pioneering masked token prediction for multi-modal generation, bridging text-to-video and action-to-video tasks, and moving towards general-purpose world model construction. The development of the Dreamer series demonstrates a clear trend: progressing from 2D to multi-dimensional video generation, transitioning from task-specific to general-purpose models, and shifting from closed, structured conditions toward more open and general natural-language-based conditions.
Diffusion-based Image Generation. The evolution of this type in autonomous driving are coalescing around two complementary thrusts: controllable generation (e.g., BEVControl (Yang et al., 2023b), DrivingDiffusion (Li et al., 2024d), and GeoDrive (Chen et al., 2025c)), which use latent diffusion to render geometry-consistent videos from rich cues such as BEV layouts, text prompts and optical flow, and high-fidelity spatio-temporal modelling, where Drive-WM (Wang et al., 2024a), Vista (Gao et al., 2024c), and LongDWM (Wang et al., 2025b) push diffusion roll-outs to longer horizons and higher spatial resolutions, markedly boosting temporal coherence and per-frame detail. The field progresses through breakthroughs in multi-modal conditions: integrating ego trajectory, human-pose, and DINO appearance tokens for behaviour-aware video synthesis (Hassan et al., 2024); translating simulator layouts into photorealistic scenes through a cascade diffusion pipeline (Zhou et al., 2024) or dual-scale controlNet-stable diffusion (Li et al., 2025c); uniting LiDAR and camera streams in a BEV latent grid for geometry-consistent cross-sensor forecasting (Zhang et al., 2024a); aligning ego–other vehicle trajectories in latent space to enable fully controllable, multi-agent video generation (Zhu et al., 2025b); and grounding multi-camera driving video generation in physics by aligning ego-world coordinates, injecting 3D flows, and box-guided occlusion reasoning (Yang et al., 2024c). Recent methods (Liang et al., 2025b; Li et al., 2025e; Jiang et al., 2025b; Ji et al., 2025; Russell et al., 2025; Wang et al., 2025a; Ma et al., 2024b; Wang et al., 2025c) focus on expanding from simple RGB image generation to more diverse and enriched paradigms, including extending conventional RGB prediction to jointly generate both images and depth (Liang et al., 2025b); shifting from layout-conditioned synthesis to fine-grained trajectory-conditioned control (Li et al., 2025e); moving beyond 2D-only conditioning toward multi-view consistent generation (Jiang et al., 2025b); transitioning from 2D layout cues to full 3D scene conditioned synthesis (Ji et al., 2025); broadening single-view synthesis into multi-view, multi-agent interactive scenarios (Russell et al., 2025); progressing from short-term clips to long-horizon video generation (Wang et al., 2025a); and further leveraging failure-case augmentation for robustness (Ma et al., 2024b); alongside the release of accident-anticipation benchmarks for risk-aware evaluation (Guan et al., 2025). Moreover, InfiniCube (Lu et al., 2025) employs world-guided video diffusion grounded on HD maps, 3D bounding boxes, and text, then lifts the generated long-horizon videos into dynamic 3D Gaussian scenes via a fast feed-forward reconstruction pipeline. This technical evolution reveals a clear trend from isolated scene generation to integrated closed-loop simulation systems, providing high-fidelity, scalable synthetic environments for autonomous driving.
Transformer-based Image Generation. Transformer-based driving imagers are evolving from narrow, single-view renderers into holistic, action-aware world models. Cross-view architectures such as HoloDrive (Wu et al., 2024b) and BEVGen (Swerdlow et al., 2024) fuse camera and LiDAR or lift BEV layouts to street-level frames, unifying 2D appearance with 3D geometry. A parallel line scales sequence modeling: DrivingWorld (Hu et al., 2024a) and GAIA-1 (Hu et al., 2023a) treat video, text and control as one token stream, enabling minute-long, controllable clips with coherent semantics. These token factories prove that unsupervised prediction alone can encode road rules and actor intent. Complementary work probes limits and remedies: DriveSim (Sreeram et al., 2024) exposes causal hallucinations in multimodal LLMs, while a DINOv2-based multi-view encoder (Popov et al., 2024) compresses synchronized front views into a latent scene grid that decodes to both RGB and BEV semantics, reducing covariate shift.
@@ 3.1.1. Image-based Generation
Table 1. Overview of summary, methods and core architectures for Image-/BEV-based generation (§3.1.1 and §3.1.2).
@@
Image-based Generation Dreamer Series DriveDreamer(Wang et al., 2024c) ECCV’24 Diffusion Model Text + Image + HDMap + Box + Actions 2D Image + Actions nuScenes(Caesar et al., 2020)
@@
DriveDreamer-2(Zhao et al., 2025b) AAAI’25 LLM + Diffusion Model Text + HDMap + Box 2D Image nuScenes(Caesar et al., 2020)
@@
DriveDreamer4D(Zhao et al., 2025a) CVPR’24 Diffusion Model + 4DGS Text + Image + Trajectory 2D Image Waymo(Sun et al., 2020)
ReconDreamer(Ni et al., 2024) CVPR’24 Diffusion Model Image + HDMap + Box 2D Image Waymo(Sun et al., 2020)
@@
WorldDreamer(Wang et al., 2024d) arXiv’24 LLM Image + Video + Text + Action 2D Image nuScenes(Caesar et al., 2020)
@@
CarDreamer(Gao et al., 2024a) IOTJ’24 World Model Backbone Image + BEV + LiDAR + Radar 2D Image + Actions CARLA(Dosovitskiy et al., 2017a)
@@
[1pt/1pt] Image-based Generation Diffusion Models BEVControl (Yang et al., 2023b) arXiv’23 Diffusion Model BEV + Image + Text 2D Image nuScenes(Caesar et al., 2020)
DrivingDiffusion(Li et al., 2024d) ECCV’24 Diffusion Model Image + Layouts + Text + Optical Flow 2D Image CARLA(Dosovitskiy et al., 2017a)
Drive-WM(Wang et al., 2024a) CVPR’24 Diffusion Model Image + Action + Box + Map + Text 2D Image nuScenes(Caesar et al., 2020) + Waymo(Sun et al., 2020)
@@
DriVerse (Li et al., 2025e) ACMMM’25 Diffusion Model Image + Trajectory 2D Image nuScenes(Caesar et al., 2020) + Waymo(Sun et al., 2020)
GEM(Hassan et al., 2024) CVPR’24 Diffusion Model Image + Trajectory + Human pose + DINO features 2D Image + Depth Self-created dataset(Hassan et al., 2024)
Vista(Gao et al., 2024c) NeurIPS’24 Diffusion Model Image + Action 2D Image Self-created dataset(Gao et al., 2024c)
SimGen (Zhou et al., 2024) NeurIPS’24 Diffusion Model Image + Depth + Segmentation + Text 2D Image DIVA (Zhou et al., 2024) + nuScenes(Caesar et al., 2020)
Delphi(Ma et al., 2024b) arXiv’24 Diffusion Model Image + BEV layout + Failure Case 2D Image nuScenes(Caesar et al., 2020)
BevWorld(Zhang et al., 2024a) arXiv’24 Diffusion Model Image + Point cloud + Action 2D Image + 3D LiDAR nuScenes(Caesar et al., 2020) + CARLA(Dosovitskiy et al., 2017b)
DrivePhysica (Yang et al., 2024c) arXiv’24 Diffusion Model Image + Poses + Box + Lane + Text 2D Image nuScenes(Caesar et al., 2020)
Imagine-2-Drive (Garg and Krishna, 2024) arXiv’24 Diffusion Model Image 2D Image CARLA(Dosovitskiy et al., 2017b)
MaskGWM(Ni et al., 2025) arXiv’25 Diffusion Model Image + Text + Action 2D Image OpenDV-2K(Yang et al., 2024b) + nuScenes(Caesar et al., 2020)
EOTWM (Zhu et al., 2025b) arXiv’25 Diffusion Model Image + Text + Trajectory 2D Image nuScenes(Caesar et al., 2020)
UniFuture (Liang et al., 2025b) arXiv’25 Diffusion Model Image + Depth 2D Image + Depth nuScenes(Caesar et al., 2020)
DiVE (Jiang et al., 2025b) arXiv’25 Diffusion Model Image + Text + Box + Road condition 2D Image nuScenes(Caesar et al., 2020)
CoGen (Ji et al., 2025) arXiv’25 Diffusion Model Image + Occupancy 2D Image nuScenes(Caesar et al., 2020)
@@
GAIA-2 (Russell et al., 2025) arXiv’25 Diffusion Model Image + Action + Text 2D Image Internal dataset(Russell et al., 2025)
MiLA (Wang et al., 2025a) arXiv’25 Diffusion Model Image + Waypoints + Text + Camera parameters 2D Image nuScenes(Caesar et al., 2020)
DiST-4D (Guo et al., 2025) ICCV’25 Diffusion Model Image + Trajectory + BEV + Pose + Box RGBD image nuScenes(Caesar et al., 2020) + Waymo(Sun et al., 2020)
SimWorld(Li et al., 2025c) arXiv’25 Diffusion Model Mask + Box + Text 2D Image AutoMine(Li et al., 2022a)
InfiniCube(Lu et al., 2025) ICCV’25 Diffusion Model + 3DGS HD Map + Bounding boxes + Text 2D Image + 3DGS Waymo(Sun et al., 2020)
STAGE(Wang et al., 2025c) arXiv’25 Diffusion Model Image + HD Map + 3D Box 2D Image + Long-horizon video nuScenes(Caesar et al., 2020)
AoTA(Guan et al., 2025) Comm. Eng.’25 Diffusion Model Image + HD Map + Layouts + Ego pose + Text 2D Image nuScenes(Caesar et al., 2020)
GeoDrive (Chen et al., 2025c) arXiv’25 Diffusion Model Image + Trajectory 2D Image nuScenes(Caesar et al., 2020)
LongDWM (Wang et al., 2025b) arXiv’25 Diffusion Model Image + Action 2D Image + Long-horizon video nuScenes(Caesar et al., 2020)
[1pt/1pt] Image-based Generation Transformers GAIA-1(Hu et al., 2023a) arXiv’23 Transformer Image + Action + Text 2D Image Internal dataset(Hu et al., 2023a)
HoloDrive(Wu et al., 2024b) arXiv’24 Transformer Image + LiDAR + Layouts + Text 2D Image + 3D LiDAR nuScenes(Caesar et al., 2020)
BEVGen(Swerdlow et al., 2024) RA-L’24 Transformer Image + BEV layout 2D Image nuScenes(Caesar et al., 2020) + Argoverse 2(Wilson et al., 2023)
DrivingWorld(Hu et al., 2024a) arXiv’24 Video GPT Image + Orientation + Location 2D Image NuPlan(Caesar et al., 2021) Internal data(Hu et al., 2024a)
DriveSim (Sreeram et al., 2024) arXiv’24 MLLMs Image + Text 2D Image Internal dataset(Sreeram et al., 2024)
@@
TMPE (Popov et al., 2024) arXiv’24 Tansformer Image + Path + Speed 2D Image + BEV CARLA (Dosovitskiy et al., 2017a)
@@
NeMo(Huang et al., 2025b) ECCV’24 Transformer Image 2D Image + LiDAR + Occupancy nuScenes(Caesar et al., 2020)
@@
FUTURIS(Karypidis et al., 2025) arXiv’25 Transformer Segmentation + Depth Segmentation + Depth Cityscapes(Cordts et al., 2016)
@@
BEV-based Generation FIERY (Hu et al., 2021b) ICCV’21 Transformer Image + Ego motion BEV nuScenes(Caesar et al., 2020)
StretchBEV(Akan and Güney, 2022) ECCV’22 LSTM Image BEV nuScenes(Caesar et al., 2020)
MILE(Hu et al., 2022b) NeurIPS’22 RNN Image + BEV segmentation + Action Image + BEV + Action CARLA(Dosovitskiy et al., 2017b)
UNO(Agro et al., 2024) CVPR’24 CNN LiDAR Occupancy + LiDAR + BEV nuScenes(Caesar et al., 2020)
PowerBEV (Li et al., 2023) IJCAI’23 CNN Image BEV Instance nuScenes(Caesar et al., 2020)